OpenStudy (anonymous):

What is the solution to the equation? 4x-263x=16-20 what does x equal

OpenStudy (anonymous):

@jacobrocket173 @Hero

OpenStudy (anonymous):

i give metals

OpenStudy (anonymous):

Medals*

OpenStudy (anonymous):

yes

pooja195 (pooja195):

4x−263x=16−20 4x+−263x=16+−20 (4x+−263x)=(16+−20)(Combine Like Terms) −259x=−4 −259x=−4 Divide both sides by -259.

OpenStudy (anonymous):

so x equal -4

pooja195 (pooja195):

Nope

OpenStudy (anonymous):

1

pooja195 (pooja195):

No

OpenStudy (anonymous):

XD

pooja195 (pooja195):

Hint : the answer is a fraction

OpenStudy (anonymous):

4x-26-3x

pooja195 (pooja195):

???

OpenStudy (anonymous):

that is what i meant

pooja195 (pooja195):

--4/--259?

pooja195 (pooja195):

^that?